Find dy if 2y3/2 + xy - x=0. Find dy if xy2 - 4x3/2 - y = 0
Answer #1

Eind dy if 243/+ny-n=0 2x3 Y dy tudy + y - 1 =0 žan du By dy + 4dy = 1-Y dn (37+4) dy = (1-y gdy dy= 11-7 dn 3yta dy ny² – 44